The manager of specialty outdoor store has gathered the following data concerning sales (in hundreds of dollars) of all the items sold by the store in the previous 2 months:

Item Type Number Sold Sales $ (00s)
Pants, jackets, and insulating layers 150 125
Hats and gloves 30 15
Socks and shoes 75 400
Skiing hardware 25 65
Climbing hardware 70 45



What is the percentage of dollar sales for any hardware?

6.9%

16.9%

10%

27.1%
Answer #1

Answer: 16.92 %

Explanation:

tem Type Number Sold Sales $ (00s)
Pants, jackets, and insulating layers 150 125
Hats and gloves 30 15
Socks and shoes 75 400
Skiing hardware 25 65
Climbing hardware 70 45
Total sum 350 650

Sum of Hardware sales= Skiing hardware+ Climbing hardware= 65+45= 110

Percent of hardware sales= sum of hardware sales/ total sales= 110/650= 0.1692 that is 16.92%

Please ask, if you have any doubts through the comment section. Do rate the answer